OpenStudy (anonymous):

Help pls. Solve. x^2/4 + 4X = 0 A. {0, –16} B. {–4, 0, 4} C. {0, –4} D. {0, 16}

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

honestly no.

OpenStudy (skullpatrol):

I this the equation you want to solve? $$\Huge \frac{x^2}{4} + 4x = 0$$

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Factor x and the constant terms

OpenStudy (skullpatrol):

When you solve equations you are looking for values of x that will make the equation a true statement, right @Homeschooler1 ?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

You could also use complete the square

OpenStudy (skullpatrol):

Let's multiply both sides by 4 to get rid or the 1/4 Ok?

OpenStudy (anonymous):

ok

OpenStudy (skullpatrol):

$$\Huge \frac{x^2}{4} + 4x = 0$$ $$\Huge 4( \frac {x^2}{4} + 4x) = 4*0$$ $$\Huge {x^2} + 16x = 0$$

OpenStudy (anonymous):

okay. Then cant we use the Zero Product Property?

OpenStudy (skullpatrol):

Yes, we will after the next step...can you tell me what that is?

OpenStudy (anonymous):

subtracting the 16x?

OpenStudy (skullpatrol):

Nope. Hint: we need to turn that equation into a product first to use the Zero product property.

OpenStudy (anonymous):

....

OpenStudy (skullpatrol):

Any other ideas? (like factoring?) :-)

OpenStudy (anonymous):

umm factoring out the equation?

OpenStudy (skullpatrol):

Yes, see the common factor of x in the two terms?

OpenStudy (skullpatrol):

$$\Huge {x^2} + 16x = 0$$

OpenStudy (anonymous):

yeah

OpenStudy (skullpatrol):

So, factor it out.

OpenStudy (skullpatrol):

$$\Huge x(x + 16) = 0$$

OpenStudy (anonymous):

ok

OpenStudy (skullpatrol):

Now, as you said, use the zero product property to find the values of x that will turn that equation into a true statement :-)

OpenStudy (anonymous):

so it would be {0, -16} ?

OpenStudy (anonymous):

Thank you for all your help!!

OpenStudy (skullpatrol):

You are correct. Thanks for asking.